OpenStudy (anonymous):

Evaluate the expression. (1p6)(12p2) a.665,280 b.132 c.5040 d.462

OpenStudy (anonymous):

what does (1p6) mean?

OpenStudy (anonymous):

i am going to assume that " 12 P 2" means "12 permute 2" i.e. \(12\times 11\)

OpenStudy (anonymous):

in the question the P is big and number are small

OpenStudy (anonymous):

yeah that is what i figured, which is why 1P 6 doesn't make any sense

OpenStudy (anonymous):



OpenStudy (anonymous):

something like that lol

OpenStudy (anonymous):

can you post it? or use the equation editor? because 1 P 6 is not right

OpenStudy (anonymous):

@satellite73 what do you mean post it?

OpenStudy (anonymous):

is it from a book? is it on line? scan or link

OpenStudy (anonymous):

it's online, I'd send the link but it's on an open exam for a class

OpenStudy (anonymous):

in "n P k" n should be greater than or equal to k

OpenStudy (anonymous):

hold on i'll take a picture an attach the file

OpenStudy (anonymous):

ok if it is an exam pretend i didn't ask you you didn't say it was an exam

OpenStudy (anonymous):

technically it's not an exam.

OpenStudy (anonymous):

that is not a 1, it is a 7 !!!

OpenStudy (anonymous):

oh then that would make sense!

OpenStudy (anonymous):

I'm sorry I couldn't tell. I still need help on this one though.

OpenStudy (anonymous):

ok well what you really need is a calculator 7 P 6 = 7!

OpenStudy (anonymous):

and 12 P 2 = 12 * 11 = 132